# WildGaussians: 3D Gaussian Splatting in the Wild

### v1

@@ 4.1 Comparison on the NeRF On-the-go Dataset
[…]
@@ 4.2 Comparision on Photo Tourism
Table 2 and Fig. 5 show results on on the challenging Photo Tourism dataset. As for the NeRF On-the-go dataset, our method shows notable improvements over all NeRF-based baselines, while enabling real-time rendering (similar to 3DGS). Compared to 3DGS, we can adeptly handle changes in appearance such as day-to-night transitions without sacrificing fine details. This shows the efficacy of our appearance modeling. Compared to the NeRF-based baseline K-Planes [7], our method offers shaper details, as can be noticed in the flowing water and text on the Trevi Fountain. Finally, compared to 3DGS [12], our method has comparable rendering speed on the NeRF On-the-go dataset, while being much faster on the Photo Tourism dataset [29]. This is caused by 3DGS [12] trying to grow unnecessary Gaussians to explain higher gradients due to the appearance variation.
@@ 4.3 Ablation Studies & Analysis
To validate the importance of each component of our method, we conduct an ablation study in Table 3, separately disabling either uncertainty or appearance modeling. Table 3 shows that without appearance modeling, performance significantly drops on the Photo Tourism dataset due to the strong appearance changes captured by the dataset. On the other dataset, which exhibit little to no illumination or other appearance changes, disabling appearance modeling only slighty improves performance (if at all). We conclude that it is safe to use appearance embeddings, even if there might not be strong appearance changes.
@@
Behavior of the appearance embedding. Fig. 7 interpolates between two appearance embeddings. The transition from the source view to the target view’s appearance is smooth, with lights gradually appearing. This demonstrates the smoothness and the continuous nature of the embedding space. Next, we further analyze the embedding space with a t-SNE [33] projection of the embeddings of training images. The t-SNE visualization in Fig. 6 reveals that the embeddings are grouped by image appearances, e.g., with night images clustering together and being separated from other images.
@@ 5 Conclusion
Limitations. While our method enables appearance modeling with real-time rendering, it is currently not able to capture highlights on objects. Additionally, although the uncertainty modeling is more robust than MSE or SSIM, it still struggles with some challenging scenarios. One way to handle this is to incorporate additional priors such as pre-trained diffusion models. We leave it as future work.
@@
We would like to thank Weining Ren for his help with the NeRF On-the-go dataset and code and Tobias Fischer and Xi Wang for fruitful discussions. This work was supported by the Czech Science Foundation (GAČR) EXPRO (grant no. 23-07973X), and by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254).
@@ A.1 Implementation & Experimental Details
NeRF On-the-go Dataset. We optimize each representation for $30k$ training steps. For our choice of learning rates, we mostly follow 3DGS [12]. We differ in the following learning rates: appearance MLP lr. of $0.0005$, uncertainty lr. of $0.001$. Gaussian embedding lr. of $0.005$. image embedding lr. of $0.001$. For the position learning rate, we exponentially decay from $1.6\times 10^{-4}$ to $1.6\times 10^{-6}$. Furthermore, we set the densification threshold at $0.0002$ and density from $500$-th iteration to 15 00015\,000-th iteration every $100$ steps. Furthermore, we reset the opacity every 3 0003\,000 steps. We do not optimize the uncertainty predictor for 500 after each opacity reset, and we do not apply the masking for the first $2000$ training steps.
Photo Tourism. We optimize each representation for $200k$ training steps. We use the following learning rates: scales lr. of $0.0005$, rotation lr. of $0.001$, appearance MLP lr. of $0.0005$, uncertainty lr. of $0.001$. Gaussian embedding lr. of $0.005$. image embedding lr. of $0.001$. For the position learning rate, we exponentially decay from $1.6\times 10^{-5}$ to $1.6\times 10^{-7}$. We set the densification threshold at $0.0002$ and density from $4000$-th iteration to 100 000100\,000-th iteration every $400$ steps. Furthermore, we reset the opacity every 3 0003\,000 steps. We do not optimize the uncertainty predictor for $1500$ after each opacity reset. We do not use the uncertainty predictor for masking for the first 35 00035\,000 steps of the training, and after that, we linearly increase its contribution for 5 0005\,000 steps.
[…]
@@
Sky Initialization. We set the scene radius $r_{s}$ to the 97% quantile of the L2 norms of the centered input 3D points. We then initialize a sphere at a distance of $10r_{s}$ from the scene center [13]. For an even distribution of points on the sphere, we utilize the Fibonacci sphere sampling algorithm [30], which arranges points in a spiral pattern using a golden ratio-based formula. We sample 100 000100\,000 points on the sphere and then project them to all training cameras, removing any points not visible from at least one camera. This set of sky points is then added to our initial point set, with their opacity initialized at 1.0, while the opacity of the rest is set at 0.1.
@@ A.2 Extended Results on NeRF On-the-go Dataset
For reference, we extend the averaged results for the NeRF On-the-go dataset [26], by giving detailed results for individual scenes. The results are presented in Table˜5.
@@ A.3 Extended Ablation Study
To further analyze the performance of the proposed contributions, we performed a detailed ablation study of both the uncertainty prediction and appearance modeling. The results are presented in Table˜4. As we can see from the table, WildGaussians appearance modeling outperforms other baselines. While VastGaussian [18] seems to work well for cases where the appearance difference is small, it fails in cases with large appearance changes and causes noticeable artifacts in the images. In the case of a simple affine color transformation (w/o Gaussian embeddings), this modeling is not powerful enough to capture local appearance changes such as lamps turning on. We can also see the effectiveness of the uncertainty modeling. However, for the Trevi Fountain, the uncertainty does not improve the performance, which is likely caused by 1) the scene has a small number of occlusions, 2) the water in the scene is often mistaken for a transient object as it is not multi-view consistent. For reference, we also included a comparison with a method trained with explicit segmentation masks obtained from the MaskRCNN predictor [9].
We also present qualitative results in Figure˜8. Notice how VastGaussians leave noticeable artifacts in the sky region in the first row. In the second row, notice how no appearance modeling or only using image embeddings cannot represent the dark sky. Similarly, not using Gaussian embeddings causes the method not to be able to represent shadows (row 2) and highlights (row 5). Disabling uncertainty leads to noticeable artifacts in row 6.

### v2

@@ 4.1 Comparison on the NeRF On-the-go Dataset
As shown in Table 1 and Fig. 4, our approach significantly outperforms both baselines, especially for scenarios with medium (15-20%) to high occlusions (30%). Compared to NeRF On-the-go [31], our method is not only 400$\times$ faster in rendering, but can more effectively remove occluders. Moreover, we can better represent distant and less-frequently seen background regions (first and third row in Fig. 4). Interestingly, 3DGS and its derivatives (Mip-Splatting, Gaussian Opacity Fields) are quite robust to scenes with low occlusion ratios, thanks to its geometry prior in the form of the initial point cloud. Nevertheless, 3DGS and its derivatives struggles to remove occlusions for high-occlusion scenes. This demonstrates the effectiveness of our uncertainty modeling strategy.
@@ 4.2 Comparision on Photo Tourism
Table 2 and Fig. 5 show results on the challenging Photo Tourism dataset. As for the NeRF On-the-go dataset, our method shows notable improvements over all NeRF-based baselines, while enabling real-time rendering (similar to 3DGS). Compared to 3DGS, we can adeptly handle changes in appearance such as day-to-night transitions without sacrificing fine details. This shows the efficacy of our appearance modeling. Compared to the NeRF-based baseline K-Planes [9], our method offers shaper details, as can be noticed in the flowing water and text on the Trevi Fountain. Compared to 3DGS [14], our method has comparable rendering speed on the NeRF On-the-go dataset, while being much faster on the Photo Tourism dataset [35]. This is caused by 3DGS [14] trying to grow unnecessary Gaussians to explain higher gradients due to the appearance variation. Finally, compared to other 3DGS-based methods [52, 5], ours achieves stronger performance while having faster inference because we can ‘bake’ the apperance-tuned spherical harmonics back into the standard 3DGS representation.
@@ 4.3 Ablation Studies & Analysis
To validate the importance of each component of our method, we conduct an ablation study in Table 3, separately disabling either uncertainty or appearance modeling. Table 3 shows that without appearance modeling, performance significantly drops on the Photo Tourism dataset due to the strong appearance changes captured by the dataset. On the NeRF On-the-go dataset, which exhibit little to no illumination or other appearance changes, disabling appearance modeling only slightly improves performance. We conclude that it is safe to use appearance embeddings, even if there might not be strong appearance changes. Similarly, disabling uncertainty modeling has little impact on datasets with less occlusions and could even make the performance slightly worse for the On-the-go low dataset, but it is required for the high-occlusion datasets (On-the-go high and Photo Tourism).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Fixed appearance multi-view consistency. We shows the multiview consistency of a fixed nighttime appearance embedding as the camera moves around the fountain.
@@
Behavior of the appearance embedding. Fig. 6 interpolates between two appearance embeddings. The transition from the source view to the target view’s appearance is smooth, with lights gradually appearing. This demonstrates the smoothness and the continuous nature of the embedding space. In Fig. 7, we interpolate between two camera poses with a fixed appearance embedding showing multiview consistency. Next, we further analyze the embedding space with a t-SNE [39] projection of the embeddings of training images. The t-SNE visualization in Fig. 8 reveals that the embeddings are grouped by image appearances, e.g., with night images clustering together and being separated from other images.
@@ 5 Conclusion
Limitations. While our method enables appearance modeling with real-time rendering, it is currently not able to capture highlights on objects. Additionally, although the uncertainty modeling is more robust than MSE or SSIM, it still struggles with some challenging scenarios. If there are not enough observations of a part of the scene, e.g., because it is occluded in nearly all training images, our approach will struggle to correctly reconstruct the region. One way to handle this is to incorporate additional priors such as pre-trained diffusion models. We leave it as future work.
@@
We would like to thank Weining Ren for his help with the NeRF On-the-go dataset and code and Tobias Fischer and Xi Wang for fruitful discussions. This work was supported by the Czech Science Foundation (GAČR) EXPRO (grant no. 23-07973X), and by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). Jonas Kulhanek acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under ELISE Grant Agreement No 951847.
@@ A.1 Implementation & Experimental Details
NeRF On-the-go Dataset. We optimize each representation for $30k$ training steps. For our choice of learning rates, we mostly follow 3DGS [14]. We differ in the following learning rates: appearance MLP lr. of $0.0005$, uncertainty lr. of $0.001$. Gaussian embedding lr. of $0.005$. image embedding lr. of $0.001$. For the position learning rate, we exponentially decay from $1.6\times 10^{-4}$ to $1.6\times 10^{-6}$. Furthermore, we set the densification threshold at $0.0002$ and density from $500$-th iteration to 15 0001500015\,00015 000-th iteration every $100$ steps. Furthermore, we reset the opacity every 3 00030003\,0003 000 steps. We do not optimize the uncertainty predictor for 500 after each opacity reset, and we do not apply the masking for the first $2000$ training steps.
Photo Tourism. We optimize each representation for $200k$ training steps. We use the following learning rates: scales lr. of $0.0005$, rotation lr. of $0.001$, appearance MLP lr. of $0.0005$, uncertainty lr. of $0.001$. Gaussian embedding lr. of $0.005$. image embedding lr. of $0.001$. For the position learning rate, we exponentially decay from $1.6\times 10^{-5}$ to $1.6\times 10^{-7}$. We set the densification threshold at $0.0002$ and density from $4000$-th iteration to 100 000100000100\,000100 000-th iteration every $400$ steps. Furthermore, we reset the opacity every 3 00030003\,0003 000 steps. We do not optimize the uncertainty predictor for $1500$ after each opacity reset. We do not use the uncertainty predictor for masking for the first 35 0003500035\,00035 000 steps of the training, and after that, we linearly increase its contribution for 5 00050005\,0005 000 steps.
Appearance MLP Priors. Considering that the appearance MLP shares weights across all Gaussians, we observed potential instability in training when randomly initializing the MLP. To mitigate this, we introduce a prior to the appearance MLP $f_{\theta}$, optimizing gradient flow during early training phases. The raw output of the last layer of the MLP, $(\hat{\beta},\hat{\gamma})$, is adjusted to obtain the affine color transformation in Eq. (5) as $\beta_{k}=0.01\hat{\beta_{k}}$ and $\gamma_{k}=0.01\hat{\gamma_{k}}+1$. This adjustment scales both the initialization of the last layer and the learning rates of the MLP by 0.01, stabilizing early-stage training such that the gradient wrt. SH coefficients is predominant.
@@
Sky Initialization. We set the scene radius $r_{s}$ to the 97% quantile of the L2 norms of the centered input 3D points. We then initialize a sphere at a distance of $10r_{s}$ from the scene center [15]. For an even distribution of points on the sphere, we utilize the Fibonacci sphere sampling algorithm [36], which arranges points in a spiral pattern using a golden ratio-based formula. We sample 100 000100000100\,000100 000 points on the sphere and then project them to all training cameras, removing any points not visible from at least one camera. This set of sky points is then added to our initial point set, with their opacity initialized at 1.0, while the opacity of the rest is set at 0.1.
@@ A.2 Extended Results on NeRF On-the-go Dataset
For reference, we extend the averaged results for the NeRF On-the-go dataset [31], by giving detailed results for individual scenes. The results are presented in Table 5.
@@
Methods were trained and evaluated on NVIDIA A100, while the rest used NVIDIA GTX 4090.
@@ A.3 Extended Ablation Study
To further analyze the performance of the proposed contributions, we performed a detailed ablation study of both the uncertainty prediction and appearance modeling. The results are presented in Table 4. As we can see from the table, WildGaussians appearance modeling outperforms other baselines. While VastGaussian [22] seems to work well for cases where the appearance difference is small, it fails in cases with large appearance changes and causes noticeable artifacts in the images. In the case of a simple affine color transformation (w/o Gaussian embeddings), this modeling is not powerful enough to capture local appearance changes such as lamps turning on. We can also see the effectiveness of the uncertainty modeling. However, for the Trevi Fountain, the uncertainty does not improve the performance, which is likely caused by 1) the scene has a small number of occlusions, 2) the water in the scene is often mistaken for a transient object as it is not multi-view consistent. For reference, we also included a comparison with a method trained with explicit segmentation masks obtained from the MaskRCNN predictor [11].
We also present qualitative results in Figure 9. Notice how VastGaussians leave noticeable artifacts in the sky region in the first row. In the second row, notice how no appearance modeling or only using image embeddings cannot represent the dark sky. Similarly, not using Gaussian embeddings causes the method not to be able to represent shadows (row 2) and highlights (row 5). Disabling uncertainty leads to noticeable artifacts in row 6.
A.4 Dataset occlusions
Refer to captionPhoto Tourism [35] (3.5%)Refer to captionRefer to captionNeRF On-the-go [31] low. (5%)Refer to captionRefer to captionNeRF On-the-go [31] high. (26%)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Occluders present in the Photo Tourism [35] and NeRF On-the-go [31] datasets.
To get some understanding of the types of occlusions present in the datasets, we visualize some images from the datasets with various amounts of occlusions in Figure 10. While for Photo Tourism [35], the occluders are humans looking into the camera in the bottom part of the images, for NeRF On-the-go [31], we have humans and objects present in various regions of the images.
A.5 Licenses
Our renderer code is based on the 3DGS codebase [14] with the modifications from Mip-Splatting [50] and Gaussian Opacity Fields [51]. They are released under “research only” licenses: https://raw.githubusercontent.com/graphdeco-inria/gaussian-splatting/refs/heads/main/LICENSE.md, https://raw.githubusercontent.com/autonomousvision/mip-splatting/refs/heads/main/LICENSE.md, and https://raw.githubusercontent.com/autonomousvision/gaussian-opacity-fields/refs/heads/main/LICENSE.md. The NeRF On-the-go dataset is licensed under the Apache 2.0 license (https://raw.githubusercontent.com/cvg/nerf-on-the-go/refs/heads/master/LICENSE). The pictures in the Photo Tourism dataset were sourced from various creators who made the images available under permissive licenses (https://creativecommons.org/share-your-work/cclicenses/). For details and the full list of creators, please refer to the original Photo Tourism paper [35].